OpenStudy (anonymous):

solve 3x=2x^2-2 by factoring

OpenStudy (anonymous):

first transpose 3x to LHS so it becomes 2x^2-3x-2=0 now it cn be represented as 2x^2-4x+x-2=0 group them 2x(x-2)+(x-2)=0 take x-2 as common (x-2)(2x+1)=0 so x=2 x=-1/2